# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Non-Hermitian skin effects (NHSEs) have recently been investigated extensively at the single-particle level. When many-body interactions become dominant, novel non-Hermitian physical phenomena can emerge. In this work, we theoretically study NHSEs controlled by dissipation and interaction. We consider a 1D zigzag Bose-Hubbard lattice, subject to magnetic flux, staggered onsite single-particle loss, and uniform onsite two-particle loss. When the two-particle loss is small, two-body bound eigenstates (i.e., doublons) are all localized at the same boundary due to the interplay of the magnetic flux and staggered single-particle loss. While, for strong two-particle loss, the localization direction of doublons is unexpectedly reversed. This is attributed to the effective strong nonreciprocal hopping of doublons induced by the virtual second-order and third-order hopping processes of particle pairs with the combination of the magnetic flux, staggered single-particle loss, and strong two-particle loss. Moreover, a two-particle gain can induce the same skin-localization and reversal features of doublons, which can be utilized to dynamically observe the NHSE and its reversal of doublons controlled by interactions. Our results open up a new avenue for exploring novel non-Hermitian phenomena in many-body systems.
@@
Introduction.—In the framework of non-Hermitian quantum mechanics Ashida et al. (2020), open quantum systems described by effective non-Hermitian Hamiltonians exhibit many striking physical phenomenon without Hermitian counterparts Peng et al. (2014a, b); Gao et al. (2015); Leykam et al. (2017); Xu et al. (2017); Gong et al. (2018); El-Ganainy et al. (2018); Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019); Zhang et al. (2018); Bliokh et al. (2019); Ge et al. (2019); Minganti et al. (2019); Zhao et al. (2019); Bliokh et al. (2019); Kawabata et al. (2019b); Borgnia et al. (2020); Minganti et al. (2020); Arkhipov et al. (2020); Okuma et al. (2020); Liu et al. (2021); Bergholtz et al. (2021); Li et al. (2022); Zhang et al. (2022a); Ren et al. (2022); Kawabata et al. (2023); Zhang et al. (2023); Li et al. (2023a); Leefmans et al. (2022); Okuma and Sato (2023); Li et al. (2024); Liu et al. (2023); Kawabata et al. (2024); Cai et al. (2024); Parto et al. (2023); Nie et al. (2023); Hu et al. (2024); Leefmans et al. (2024). One of the fascinating properties of non-Hermitian systems is the localization of bulk modes at the boundaries Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019), dubbed the non-Hermitian skin effect (NHSE). The NHSE originates from the intrinsic point-gap topology of complex eigenenergies Okuma et al. (2020). Recently, many efforts have been devoted to exploring unique consequences of NHSE without many-body interactions. For instance, it leads to the breakdown of conventional Bloch band theory Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020), nonunitary scaling of non-Hermitian localization Kawabata and Ryu (2021), and entanglement phase transitions Kawabata et al. (2023).
@@
When many-body interactions are introduced into non-Hermitian systems, unusual physical properties can emerge Ashida et al. (2017); Nakagawa et al. (2018); Yamamoto et al. (2019); Nakagawa et al. (2020); Yoshida and Hatsugai (2021); Zhang et al. (2022b); Dóra and Moca (2022); Kawabata et al. (2022); Liu et al. (2020); Yoshida and Hatsugai (2022); Shen and Lee (2022); Faugno and Ozawa (2022); Li et al. (2023b); Longhi (2023); Qin and Li (2024); Yoshida et al. (2023); Qin et al. (2024); Brighi and Nunnenkamp (2024); Kim et al. (2024). Especially, for interacting systems in the presence of nonreciprocal hopping, many-body interactions have the potential to induce NHSE Faugno and Ozawa (2022), occupation-dependent NHSE Qin and Li (2024), non-Hermitian Mott skin effect Yoshida et al. (2023), dynamical suppression of NHSE Qin et al. (2024), and so on. However, it is challenging requirement of nonreciprocal hopping in achieving the NHSE for many experimental platforms. A more operational approach is to utilize onsite loss or gain Li et al. (2020); Yi and Yang (2020); Li et al. (2022); Liang et al. (2022); Cai et al. (2024). Therefore, a natural open question arises: can the onsite loss or gain be utilized to induce and control NHSEs in the interacting system?
@@
In this work, we study the skin-mode localization in dissipative and interacting systems without nonreciprocal hopping. Specifically, we consider a 1D zigzag lattice in the presence of a magnetic flux piercing through each triangular plaquette and a Bose-Hubbard interaction. For strong interaction, bound pairs, so-called doublon, are formed, where two bosons occupy the same site. By introducing the imbalanced onsite single-particle loss at the odd and even sites and uniform onsite two-particle loss, two branches of doublon bands form. For weak two-particle loss, the states from two doublon bands are all localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Most remarkably, the states from one branch of the doublon bands are unexpectedly reversely localized at the other boundary for strong two-particle loss. The underlying hidden physical mechanism is revealed by utilizing third-order quasidegenerate perturbation theory, where an effectively strong nonreciprocal hopping contributes from the virtual second-order and third-order hopping processes of particle pairs with the combination of the magnetic flux, staggered single-particle loss, and strong two-particle loss. We finally discuss the same reversal of the NHSE induced by adding two-particle gain. Moreover, with two-particle gain, the initial two-particle excitation can remain strongly bound during the dynamical evolution, which can be utilized to observe dynamical skin-mode localization and its reversal of doublons controlled by the interactions.
@@
$-\sum_{j=1}^{L}t\left(\hat{a}^{\dagger}_{j+1}\hat{a}_{j}+\text{H.c.}\right)-U\sum_{j=1}^{L}\hat{a}^{\dagger}_{j}\hat{a}^{\dagger}_{j}\hat{a}_{j}\hat{a}_{j},$ (1)
@@
Figure 1: Tight-binding representation of a two-leg zigzag chain in the presence of onsite interaction $U$, imbalanced onsite single-particle loss (denoted by purple wavy lines) with decay rate $\gamma\pm\delta$ on two legs, and uniform onsite two-particle loss (denoted by yellow wavy lines) with decay rate $\kappa$. A synthetic magnetic flux $\phi$ piercing through each triangular plaquette is introduced. The black solid and red dashed lines represent the nearest-neighbor and next-nearest-neighbor hopping with strengths $t$ and $t^{\prime}$.
We consider this open quantum system, where the zigzag lattice is subjected to the imbalanced onsite single-particle loss between two legs, and uniform onsite two-particle inelastic collision (see Fig. 1). Under the Markovian approximation, the dissipative dynamics of the system is governed by the Lindblad master equation Scully and Zubairy (1997); Breuer and Petruccione (2007); Lidar (2020)
@@
where $\hat{\rho}$ is the system density matrix, and the Lindblad superoperator $\mathcal{D}[\hat{\mathcal{L}}]\hat{\rho}=\hat{\mathcal{L}}\hat{\rho}\hat{\mathcal{L}}^{\dagger}-\{\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}},~{}\hat{\rho}\}/2$ represents bosonic dissipation. The imbalanced single-particle losses along the two legs are denoted by $\gamma\pm\delta$, and $\kappa$ is the two-particle decay rate.
Quantum-trajectory theory states that the dynamics in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) can be decomposed into nonunitary evolution and stochastic quantum jumps Meystre and Sargent (2007). By projecting out the quantum jumps by continuously monitoring the particle number Ashida et al. (2016); Nakagawa et al. (2020), the system is governed by the effective non-Hermitian Hamiltonian
@@
where $\hat{n}_{j}=\hat{a}_{j}^{\dagger}\hat{a}_{j}$ is the density operator. The Hamiltonian ^$\hat{\mathcal{H}}_{\text{nH}}$ conserves the $U(1)$ symmetry, i.e., ^$\hat{\mathcal{H}}_{\text{nH}}$ commutes with the total particle number operator [^$[\hat{\mathcal{H}}_{\text{nH}},~{}\sum_{j}\hat{n}_{j}]=0$. Thus, we can study the arbitrary subspace of the particle number in the Fock basis. In this work, we consider the double-excitation subspace and assume $\phi=\pi/2$.
Figure 2: (a) Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) in the presence of single-particle dissipation for $\kappa=0$. The discrete doublon bands on the far left are well separated from the continuum of scattering states. The zoom‐in view of the two doublon bands in (a) is shown in (b), where the magenta star and diamond denote the two branches of the doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, $\delta=2$, and $L=60$.
Single-Particle Dissipation Induced NHSE.—We first study the NHSE in the presence of only the single-particle dissipation with $\kappa=0$. We plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under open boundary conditions (OBCs, blue dots) and periodic boundary conditions (PBCs, red dots) in Fig. 2(a). In the double-excitation subspace, the eigenspectrum is divided into continuous scattering states and discrete doublon bands. The scattering states consist of superpositions of two-particle states residing on different sites, while the doublon bands are composed of bound-particle pairs residing on the same site, with their eigenenergies well separated from the continuum of scattering states [see Fig. 2(a)]. In this work, we focus on two doublon bands, of which the zoom-in view on the far left of Fig. 2(a) is shown in Fig. 2(b). In the complex plane, each doublon band, indicated by the magenta star and diamond, forms a point gap (red dots) under PBCs. Each of these point gaps encircles the respective doublon band (blue dots) under OBCs, indicating usually the occurrence of NHSE. In order to demonstrate this indeed the case, we calculate the particle density as
@@
where |$\ket{\psi_{m}}_{R}$ ($m=1,2,3,\cdots$) is the $m$th right eigenvector of the two-body non-Hermitian Hamiltonian with ^$\hat{\mathcal{H}}_{\text{nH}}\ket{\psi_{m}}_{R}=E_{m}\ket{\psi_{m}}_{R}$.
Figure 3: Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) subjected to single- and two-particle dissipation for $\kappa=16$ and $\delta=2$. The zoom‐in view of the two doublon bands on the far left of (a) is shown in (b), where the magenta star and diamond denote two branches of doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, and $L=60$.
The site-resolved particle densities of two doublon bands are shown in Figs. 2(c,d), respectively, where red (blue) curves indicate state distributions at odd (even) sites. All the doublons are localized at the right boundary, indicating the NHSE. Moreover, the states from two branches of doublon bands occupy different legs of the zigzag lattice, where the states from the doublon band with larger (smaller) loss reside mainly the odd (even) sites [see Fig. 2(c,d)]. In order to reveal the hidden reasons for the appearance of the doublon skin-mode localization caused by the imbalanced onsite single-particle loss, we derive its effective Hamiltonian (see details in Supplementary Materials (SM) in Ref. Non ). The NHSE of doublon bands can be attributed to the interplay of the effective magnetic flux of doublons, threading through each triangular plaquette, and the renormalized imbalanced loss between its two legs. The non-zero flux leads to opposite particle transports along the odd and even sites Li et al. (2022); Cai et al. (2019). Due to the larger loss at the odd sites, the particle transport towards the right boundary are favored, and the backflow is suppressed. This leads to the skin-mode localization at the right boundary. Note that two isolated states of doublons, lying outside the point gaps [Fig. 2(b)], are localized at either the left or right boundary. There are two-particle Tamm-Shockley edge states Tamm (1932); Shockley (1939), resulting from the different renormalization of the system parameters at the edges and bulk (see details in Ref. Non ), which have no topological origin.
@@
Two-Particle Dissipation Controlled NHSE.—We include the effects of two-particle loss, and assume the two-particle loss much larger than the single-particle one with $\kappa=16$ and $\delta=2$, and plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) in Fig. 3(a). The zoom-in view of doublon bands is shown in Fig. 3(b). The eigenenergies of OBCs are enclosed by the point-gap loop, indicating the appearance of NHSE. In contrast to the case of single-particle dissipation, the two-particle loss reverses the localization direction of doublons, as shown in Fig. 3(c,d). The states from the doublon band with the larger loss occupy the odd sites and are still localized at the right boundary [see Fig. 3(c)]. However, the states from the other doublon band with the smaller loss are unexpectedly localized at the left boundary [see Fig. 3(d)], where they occupy the even sites. In addition, the winding number of the doublon band with the smaller loss becomes $\mathcal{W}=-2$, indicating also the reversal of NHSE. Furthermore, the localization direction of the doublon skin modes can be tuned by the interaction, single-particle loss or two-particle dissipation. For instance, by fixing the two-particle loss $\kappa$, but increasing the single-particle dissipation $\delta$, we observe once more the reversal of the localization direction of the doublon band Non .
In order to understand the underlying hidden mechanism of the unexpected reversal of the localization direction of the doublon skin modes, we derive the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ describing the doublon bands in the strong-interaction limit with $U\gg t,t^{\prime}$. In this case, the doublons are tightly bound, where their two-particle components reside on the same site. Moreover, the doublon states are energetically well-separated from two-particle scattering states. Therefore, we can split the Hamiltonian ^$\hat{\mathcal{H}}_{\text{nH}}$ into its dominant term, with ^$\hat{\mathcal{H}}_{0}=-(U+i\kappa/2)\sum_{j}\hat{n}_{j}(\hat{n}_{j}-1)-\frac{i}{2}\sum_{j}\left[\gamma-(-1)^{j}\delta\right]\hat{n}_{j}$, and treat the remaining terms as the perturbation. Based on third-order quasidegenerate perturbation theory Non , we obtain the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ of doublons as
@@
According to the above doublon’s effective Hamiltonian in the 1D zigzag lattice, the magnetic flux piercing through each triangular plaquette [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)] and the imbalanced complex onsite potential $U(\pm\delta)$ [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)]) at odd and even sites are renormalized. Most remarkably, along each leg of the zigzag lattice, the virtual second-order hopping process of particle pairs contributes to an effective phase $2\phi$, and third-order process contributes to $\phi$ in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) (see details in Ref. Non ). Therefore, the effective next-nearest neighbor hopping of doublons becomes nonreciprocal, i.e., |$\absolutevalue{J_{i,\text{R}}}\neq\absolutevalue{J_{i,\text{L}}}$ | ($i=1,2$) in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects), due to the interplay of two-particle dissipation and magnetic flux. Moreover, the nonreciprocal-hopping direction can be different along two legs, determined by the interaction strength $U$, two-particle loss $\kappa$, and imbalanced single-particle loss $\delta$.
@@
According to the above discussion, in the presence of two-particle dissipation, the skin-mode localization direction of doublons is determined by the competition between the effective nonreciprocal hopping, induced by the two-particle dissipation and magnetic flux, and the directional hopping due to the interplay of magnetic-flux induced chiral transport and imbalanced loss along two legs. This contrasts with the single-particle case, where the NHSE results from the interplay of the magnetic-flux-induced chiral transport and the imbalanced onsite loss along two legs of the zigzag lattice structure in Fig. 1. By fixing $U=6$ and $\delta=2$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}<\absolutevalue{J_{2,\text{L}}}$ | for $\kappa>\delta$. If the two-particle dissipation $\kappa$ is comparable with the interaction strength $U$, all the states from two doublon bands are localized at the right boundary (see discussions with $\kappa=7$ in Ref. Non ). This is mainly attributed to the directional hopping due to the interplay of the magnetic-flux-induced chiral transport and imbalanced loss along two legs of doublons [see Eqs. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) and (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)] in spite of the slightly larger hopping strength towards the left boundary along the even sites. However, as the two-particle dissipation strength $\kappa$ is further increased, the asymmetrical hopping towards the left boundary becomes much stronger than hopping towards the right one along the even sites. This finally reverses the localization direction of the states from the doublon band with the smaller loss for $\kappa\gg U$, where they are localized at the left boundary along the even sites [see Fig. 3(d)]. Furthermore, if $\delta$ is increased and becomes comparable with $\kappa$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}>\absolutevalue{J_{2,\text{L}}}$ |. This causes once more the reversal of NHSE along the even sites [see Fig. 3(h)].

### v2

@@ Abstract
Non-Hermitian skin effects (NHSEs) have recently been investigated extensively at the single-particle level. When many-body interactions become dominant, novel non-Hermitian physical phenomena can emerge. In this work, we theoretically study NHSEs controlled by dissipation and interaction. We consider a 1D zigzag Bose-Hubbard lattice, subject to magnetic flux, staggered onsite single-particle loss, and uniform onsite two-particle loss. When the two-particle loss is small, two-body bound eigenstates (i.e., doublons) are all localized at the same boundary due to the interplay of the magnetic flux and staggered single-particle loss. While, for strong two-particle loss, the localization direction of doublons is unexpectedly reversed. This is attributed to the effective strong nonreciprocal hopping of doublons contributing from the virtual second-order and third-order hopping processes of particle pairs in combination with the magnetic flux, the strong two-particle loss, and the many-body interaction. Moreover, a two-particle gain can induce the same skin-localization of doublons, which can be utilized to dynamically observe the NHSE and its reversal of doublons controlled by interactions. Our results open up a new avenue for exploring novel non-Hermitian phenomena in many-body systems.
@@
Introduction.—In the framework of non-Hermitian quantum mechanics Ashida et al. (2020), open quantum systems described by effective non-Hermitian Hamiltonians exhibit many striking physical phenomenon without Hermitian counterparts Peng et al. (2014a, b); Gao et al. (2015); Leykam et al. (2017); Xu et al. (2017); Gong et al. (2018); El-Ganainy et al. (2018); Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019); Zhang et al. (2018); Bliokh et al. (2019); Ge et al. (2019); Minganti et al. (2019); Zhao et al. (2019); Bliokh et al. (2019); Kawabata et al. (2019b); Borgnia et al. (2020); Minganti et al. (2020); Arkhipov et al. (2020); Okuma et al. (2020); Liu et al. (2021); Bergholtz et al. (2021); Li et al. (2022); Zhang et al. (2022a); Li and Xu (2022); Ren et al. (2022); Kawabata et al. (2023); Zhang et al. (2023); Li et al. (2023a); Leefmans et al. (2022); Okuma and Sato (2023); Li et al. (2024); Liu et al. (2023); Kawabata et al. (2024); Manna and Roy (2023); Cai et al. (2024); Parto et al. (2023); Nie et al. (2023); Wang et al. (2024); Hu et al. (2024); Leefmans et al. (2024). One of the fascinating properties of non-Hermitian systems is the localization of bulk modes at the boundaries Yao and Wang (2018); Zhang et al. (2020); Yokomizo and Murakami (2019); Yao et al. (2018); Kunst et al. (2018); Liu et al. (2019); Song et al. (2019); Lee et al. (2019); Kawabata et al. (2019a); Lee et al. (2019), dubbed the non-Hermitian skin effect (NHSE). The NHSE originates from the intrinsic point-gap topology of complex eigenenergies Okuma et al. (2020). Recently, many efforts have been devoted to exploring unique consequences of NHSE without many-body interactions, e.g., breakdown of conventional Bloch band theory Yao and Wang (2018); Yokomizo and Murakami (2019); Zhang et al. (2020) and entanglement phase transitions Kawabata et al. (2023).
@@
When many-body interactions are introduced into non-Hermitian systems, unusual physical properties can emerge Ashida et al. (2017); Nakagawa et al. (2018); Yamamoto et al. (2019); Nakagawa et al. (2020); Yoshida and Hatsugai (2021); Zhang et al. (2022b); Dóra and Moca (2022); Kawabata et al. (2022); Liu et al. (2020); Yoshida and Hatsugai (2022); Shen and Lee (2022); Faugno and Ozawa (2022); Li et al. (2023b); Longhi (2023); Mao et al. (2023); Qin and Li (2024); Yoshida et al. (2024); Qin et al. (2024); Brighi and Nunnenkamp (2024); Kim et al. (2024); Begg and Hanai (2024); Mao et al. (2024). Especially, with nonreciprocal hopping, many-body interactions have the potential to induce NHSE Faugno and Ozawa (2022), occupation-dependent NHSE Qin and Li (2024), non-Hermitian Mott skin effect Yoshida et al. (2024), dynamical suppression of NHSE Qin et al. (2024), and so on. However, it is challenging to implement the nonreciprocal hopping in order to achieve the NHSE for many experimental platforms. A more operational approach is to utilize onsite loss or gain Li et al. (2020); Yi and Yang (2020); Li et al. (2022); Liang et al. (2022); Cai et al. (2024). Therefore, a natural open question arises: can the onsite loss or gain be utilized to induce and even control NHSEs in the interacting non-Hermitian system?
@@
In this work, we study the NHSE in dissipative and interacting systems without nonreciprocal hopping. Specifically, we consider a 1D zigzag Bose-Hubbard lattice in the presence of a magnetic flux piercing through each triangular plaquette. By introducing the imbalanced onsite single-particle loss at the odd and even sites and uniform onsite two-particle loss, two branches of doublon (i.e., a bound pair at the same site) bands form. For weak two-particle loss, the states from two doublon bands are all localized at the same boundary due to the interplay of the magnetic-flux-induced chiral transport and staggered single-particle loss. Most remarkably, the states from one branch of the doublon bands are unexpectedly reversely localized at the other boundary for strong two-particle loss. The underlying hidden physical mechanism is revealed: there is an effectively strong nonreciprocal hopping contributing from the virtual second-order and third-order hopping processes of particle pairs in combination with the magnetic flux, the strong two-particle loss, and the interaction. We finally discuss the same reversal of the NHSE induced by adding two-particle gain. Moreover, with two-particle gain, the initial two-particle excitation can remain strongly bound during the dynamical evolution, which can be utilized to observe dynamical NHSE and its reversal of doublons controlled by the interactions.
@@
$-\sum_{j=1}^{L-1}t\left(\hat{a}^{\dagger}_{j+1}\hat{a}_{j}+\text{H.c.}\right)-U\sum_{j=1}^{L}\hat{a}^{\dagger}_{j}\hat{a}^{\dagger}_{j}\hat{a}_{j}\hat{a}_{j},$ (1)
@@
Figure 1: Schematic showing a two-leg zigzag chain in the presence of onsite interaction $U$, imbalanced onsite single-particle loss (denoted by purple wavy lines) with decay rate $\gamma\pm\delta$ on two legs, and uniform onsite two-particle loss (denoted by yellow wavy lines) with decay rate $\kappa$. A synthetic magnetic flux $\phi$ piercing through each triangular plaquette is introduced. The black solid and red dashed lines represent the nearest-neighbor and next-nearest-neighbor hopping with strengths $t$ and $t^{\prime}$.
We suppose that the lattice undergoes the imbalanced onsite single-particle loss $\gamma\pm\delta$ between two legs, and uniform onsite two-particle loss $\kappa$ (see Fig. 1). The system dynamics is governed by the Lindblad master equation Scully and Zubairy (1997); Breuer and Petruccione (2007); Lidar (2020)
@@
where $\hat{\rho}$ is the system density matrix, and the Lindblad superoperator $\mathcal{D}[\hat{\mathcal{L}}]\hat{\rho}=\hat{\mathcal{L}}\hat{\rho}\hat{\mathcal{L}}^{\dagger}-\{\hat{\mathcal{L}}^{\dagger}\hat{\mathcal{L}},~{}\hat{\rho}\}/2$ represents bosonic dissipation.
By continuously monitoring particle number Ashida et al. (2016), the dissipative system in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) is described by the non-Hermitian Hamiltonian
@@
where $\hat{n}_{j}=\hat{a}_{j}^{\dagger}\hat{a}_{j}$ is density operator. Due to [^$[\hat{\mathcal{H}}_{\text{nH}},~{}\sum_{j}\hat{n}_{j}]=0$, we can study the arbitrary subspace of the particle number in the Fock basis. Here, we consider the double-excitation subspace and assume $\phi=\pi/2$.
Figure 2: (a) Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) for $\kappa=0$ and $\delta=2$. The zoom‐in view of the two doublon bands on the far left of (a) is shown in (b), where the magenta star and diamond denote the two branches of the doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, $\delta=2$, and $L=60$.
Single-Particle Dissipation Induced NHSE.—We first study the NHSE in the presence of only the single-particle dissipation with $\kappa=0$. We plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under open boundary conditions (OBCs, blue dots) and periodic boundary conditions (PBCs, red dots) in Fig. 2(a). The eigenspectrum is divided into continuous scattering states and discrete doublon bands. The scattering states consist of superpositions of two-particle states residing on different sites, while the doublon bands are composed of bound-particle pairs residing on the same site, with their eigenenergies well separated from the continuum of scattering states [see Fig. 2(a)]. In this work, we focus on two doublon bands, of which the zoom-in view on the far left of Fig. 2(a) is shown in Fig. 2(b). In the complex plane, each doublon band, indicated by the magenta star and diamond, forms a point gap (red dots) under PBCs. Each of these point gaps encircles the respective doublon band (blue dots) under OBCs, indicating usually the occurrence of NHSE. In order to demonstrate this indeed the case, we calculate the particle density as
@@
where |$\ket{\psi_{m}}_{R}$ ($m=1,2,3,\cdots$) is the $m$th right eigenvector with ^$\hat{\mathcal{H}}_{\text{nH}}\ket{\psi_{m}}_{R}=E_{m}\ket{\psi_{m}}_{R}$.
Figure 3: Complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) for $\kappa=16$ and $\delta=2$. The zoom‐in view of the two doublon bands on the far left of (a) is shown in (b), where the magenta star and diamond denote two branches of doublon bands. The corresponding particle densities $\langle\hat{n}_{j}\rangle$ are shown in (c) and (d) under OBCs, where the red (blue) curves indicate the state distributions at the odd (even) sites. Other parameters are: $U=6$, $t=2$, $t^{\prime}=1$, $\gamma=4$, and $L=60$.
The site-resolved particle densities of two doublon bands are shown in Figs. 2(c,d), respectively, where red (blue) curves indicate state distributions at odd (even) sites. All the doublons are localized at the right boundary, indicating the NHSE. Moreover, the states from two branches of doublon bands occupy different legs of the zigzag lattice, where the states from the doublon band with larger (smaller) loss reside mainly the odd (even) sites [see Fig. 2(c,d)]. For the single-particle loss, the NHSE of doublons is attributed to the interplay of the effective magnetic flux of doublons, and the renormalized imbalanced loss between its two legs (see the effective Hamiltonian of doublons in Supplementary Materials (SM) in Ref. Non ). The non-zero flux leads to opposite particle transports along the two legs Li et al. (2022); Cai et al. (2019). Due to the larger loss at the odd sites, the particle transport towards the right boundary are favored, and the backflow is suppressed. This leads to the skin-mode localization at the right boundary. Note that two isolated states of doublons, lying outside the point gaps [Fig. 2(b)], are localized at either the left or right boundary. There are two-particle Tamm-Shockley edge states Tamm (1932); Shockley (1939), resulting from different effective onsite potential at the edges and bulk (see details in Ref. Non ), which have no topological origin.
@@
Two-Particle Dissipation Controlled NHSE.—We include the effects of the strong two-particle loss with $\kappa=16$ and $\delta=2$, and plot the complex eigenenergies of ^$\hat{\mathcal{H}}_{\text{nH}}$ under OBCs (blue dots) and PBCs (red dots) in Fig. 3(a). The zoom-in view of doublon bands is shown in Fig. 3(b). The eigenenergies of OBCs are enclosed by the point-gap loop, indicating the appearance of NHSE. In contrast to the case of single-particle dissipation, the two-particle loss reverses the localization direction of doublons, as shown in Fig. 3(c,d). The states from the doublon band with the larger loss occupy the odd sites and are still localized at the right boundary [see Fig. 3(c)]. However, the states from the other doublon band with the smaller loss are unexpectedly localized at the left boundary [see Fig. 3(d)], where they occupy the even sites. In addition, the winding number of the doublon band with the smaller loss becomes $\mathcal{W}=-2$, indicating also the reversal of NHSE. Furthermore, the localization direction of the doublons can be tuned by the interaction, single-particle loss or two-particle dissipation. For instance, by fixing the two-particle loss $\kappa$, but increasing the single-particle dissipation $\delta$, we observe once more the reversal of the localization direction of the doublon band Non .
In order to understand the underlying hidden mechanism of the unexpected reversal of the localization direction of the doublons, we derive the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ describing the doublon bands in the strong-interaction limit with $U\gg t,t^{\prime}$. In this case, the doublons are tightly bound, and their bands are energetically well-separated from two-particle scattering states. Therefore, we can split the Hamiltonian ^$\hat{\mathcal{H}}_{\text{nH}}$ into its dominant term, with ^$\hat{\mathcal{H}}_{0}=-(U+i\kappa/2)\sum_{j}\hat{n}_{j}(\hat{n}_{j}-1)-\frac{i}{2}\sum_{j}\left[\gamma-(-1)^{j}\delta\right]\hat{n}_{j}$, and treat the remaining terms as the perturbation. Based on third-order quasidegenerate perturbation theory Non , we obtain the effective Hamiltonian ^$\hat{\mathcal{H}}_{\text{eff}}$ of doublons as
@@
As shown by the above effective Hamiltonian, the zigzag lattice of doublons is pierced through the magnetic flux [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)], where there exists the imbalanced complex onsite potential $U(\pm\delta)$ along two legs [see Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)]). Most remarkably, along each leg of the lattice, the virtual second-order hopping process of particle pairs contributes to an effective phase $2\phi$, and third-order process contributes to $\phi$ in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects) (see details in Ref. Non ). Therefore, the effective next-nearest neighbor hopping of doublons becomes nonreciprocal, i.e., |$\absolutevalue{J_{i,\text{R}}}\neq\absolutevalue{J_{i,\text{L}}}$ | ($i=1,2$) in Eq. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects), due to the interplay of the strong two-particle dissipation and the magnetic flux in the presence of the many-body interactions. Moreover, the nonreciprocal-hopping direction can be different along two legs, determined by the interaction $U$, two-particle loss $\kappa$, and imbalanced single-particle loss $\delta$.
@@
As discussed above, with the two-particle loss, the skin-mode localization direction of doublons is determined by the competition between the effective nonreciprocal hopping, induced by the two-particle dissipation and magnetic flux, and the directional hopping due to the interplay of magnetic-flux induced chiral transport and imbalanced loss along two legs. This contrasts with the single-particle case, where the NHSE results from the interplay of the magnetic-flux-induced chiral transport and the imbalanced onsite loss along two legs of the zigzag lattice structure in Fig. 1. By fixing $U=6$ and $\delta=2$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}<\absolutevalue{J_{2,\text{L}}}$ | for $\kappa>\delta$. If the two-particle dissipation $\kappa$ is comparable with the interaction strength $U$, all the states from two doublon bands are localized at the right boundary (see discussions with $\kappa=7$ in Ref. Non ). This is mainly attributed to the directional hopping due to the interplay of the magnetic-flux-induced chiral transport and imbalanced loss along two legs of doublons [see Eqs. (Dissipation and Interaction-Controlled Non-Hermitian Skin Effects, Dissipation and Interaction-Controlled Non-Hermitian Skin Effects)] in spite of the slightly larger hopping strength towards the left boundary along the even sites. However, as the two-particle dissipation strength $\kappa$ is further increased, the asymmetrical hopping towards the left boundary becomes much stronger than hopping towards the right one along the even sites. This finally reverses the localization direction of the states from the doublon band with the smaller loss for $\kappa\gg U$, where they are localized at the left boundary along the even sites [see Fig. 3(d)]. Furthermore, if $\delta$ is increased and becomes comparable with $\kappa$, we have |$\absolutevalue{J_{1,\text{R}}}>\absolutevalue{J_{1,\text{L}}}$ | and |$\absolutevalue{J_{2,\text{R}}}>\absolutevalue{J_{2,\text{L}}}$ |. This causes once more the reversal of NHSE along the even sites [see Fig. 3(h)].